OpenStudy (anonymous):

A bank in Houston loans $30,000 to a business at 14.5% annual simple interest for 21 months. How much interest will the bank earn? Can anyone help explain to me how to solve this, please?

OpenStudy (anonymous):

21 months = 1 and 3/4 years. So compute interest for 1.75 years based on 14.5% a year.

OpenStudy (anonymous):

Do I use the A = P(1 + r/n)^nt formula?

OpenStudy (anonymous):

NO, thats for inteest compounded annually..Your problem is simple interest

OpenStudy (anonymous):

Oh

OpenStudy (anonymous):

Watch:

OpenStudy (anonymous):

For one year, interest of 14.5% of $30,000 = $4,350.00

OpenStudy (anonymous):

That is just .145 multiplied by 30,000

OpenStudy (anonymous):

so we want 1 3/4 years...so it is another 3/4 of $30,000 = $3,262.50

OpenStudy (anonymous):

Total interest paid after 21 months (1 3/4 years) is $4,350 + $3,262.50 = $7,612.50

OpenStudy (anonymous):

I hope I didnt make an error.

OpenStudy (anonymous):

no, I am correct.

OpenStudy (anonymous):

Ok I think i get it.

OpenStudy (anonymous):

Great.